ON THE RENNER-TELLER EFFECT AND BARRIERS TO LINEARITY AND DISSOCIATION IN HCF (A^{1}A′′)

H. FAN, I. IONESCU, \underline{C. ANNESLEY}, AND S. A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53233; J. XIN, Department of Physics and Engineering Technologies, Bloomsburg University, Bloomsburg, PA 17815.

To further investigate the Renner-Teller effect and excited state barriers to linearity and dissociation in the simplest singlet carbene, HCF, we measured fluorescence excitation spectra and lifetimes of the pure bending transitions 2_{0}^{n} with n = 0-7 and the combination bands 1_{0}^{1}2_{0}^{n} with n = 1-6 and 2_{0}^{n}3_{0}^{1} with n = 0-3 in the HCF A^{1}A′′-X^{1}A′ system. The spectra were measured under jet-cooled conditions using a pulsed discharge source, and rotationally analyzed to yield precise values for the band origins and rotational constants. The derived A^{1}A′′ state parameters are in excellent agreement with the predictions of ab initio electronic structure theory. The approach to linearity is evidenced in a sharp increase in the A rotational constant, a minimum in the bending vibrational intervals, and a pronounced fluorescence lifetime lengthening for levels with K′_{a} > 0. A fit of the vibrational intervals for the pure bending levels yields a barrier to linearity of 6300 cm^{−1} above the vibrationless level. Our observation of the K′_{a} = 1 level of (1,6,0) places a lower limit on the barrier to dissociation of 8555 cm^{−1} above the vibrationless level.